Q#1)price of 3 football and 5 tennis ball is R.s305.the price of 2 tenis ball and 5 football is R.s. 350.find price of one fòotball.
Answer by Boreal(15235)

3F+5T=305
5F+2T=350
multiply top by 5 and the bottom by -3
15F+25T=1525
-15F-6T==-1050
19T=475
T=25
substitute into first
3F+125=305
3F=180
F=$60 ANSWER.
check in second
300+50=350.